# Aligning Visual Contrastive learning models via Preference Optimization

### v1

@@ Appendix B Proofs
$=\mathbb{E}_{(x_{1},y_{1},y_{2})\sim\mathcal{D}}\left[\left(\log\left(\frac{x_{1}}{x_{2}}\right)-\frac{\beta^{-1}}{2}\right)^{2}\right]$ (28)
@@
$=2\left(\log\left(\frac{x_{1}}{x_{2}}\right)-\frac{\beta^{-1}}{2}\right)\cdot\frac{1}{x_{1}}$
@@ B.3 The Gradient of the DPO and IPO Objectives
Appendix C What features does our method learn
In Figure 6, we showcase how applying our method, forces the model to focus on the suitable features of the data rather than different sorts of attacks such as typographic.
a photo of a dumpling.
a photo of a clam chowder.
ProbabilityInverse (typo.)Inverse (clean)Pretrained (typo)Pretrained (clean)Finetuned (typo.)Finetuned (clean)
Figure 6: Typographic attack assessment. (Left) Saliency maps of vanilla CLIP and our fine-tuned version on typographic images. Each column represents a sample consisting of a typographic image, a saliency map for the prediction of vanilla CLIP, and a saliency map for the prediction of our fine-tuned version. As seen, vanilla CLIP focuses on unrelevant parts of images, especially typo texts, which results in false predictions. On the other hand, our model is capable of ignoring such fake text and focusing on the visual concept of the image, which results in true prediction. (Center) Image after typographic attack with misleading text (”clam chowder”). (Right) Model predictions: The finetuned model remains robust, correctly classifying the attacked image, while the original model misclassifies it. The inverse model confidently selects the typographic label but preserves classification accuracy.
@@ Appendix D Latent visualization using VQGAN
In this section, we further investigate our fine-tuned CLIP on typographical attack datasets. We use VQGAN-CLIP (Crowson et al., 2022) as the generative model which use CLIP in its backbone. The results are shown in figure 7. As can be seen, our fine-tuned CLIP completely ignores text in its image encoder representation and only focus on visual concepts which increase it reliabity and trustworthiness.
@@
Figure 7: Retrieved images using VQGAN-CLIP (Crowson et al., 2022). The captions for image generation is ”focus”, ”Love”, ”Male police”, ”Time”. As can be seen, vanilla CLIP put the texts on image too, but our model fine-tuned on typographical attack dataset, SUN, ignores such image texts and generated only visual concepts.
Appendix E Datasets
@@
Tabel 3 shows the statistics of the datasets we used in this paper. To evaluate the accuracy of our method on both normal classification and typographic classification (results on table 1), we consider $9$ datasets: ImageNet-100 (Deng et al., 2009), Caltech101 (Fei-Fei et al., 2004), OxfordPets (Parkhi et al., 2012), StanfordCars (Krause et al., 2013), Flowers102 (Nilsback & Zisserman, 2008), FGVCAirCrafts (Maji et al., 2013), DTD (Cimpoi et al., 2014), SUN397 (Xiao et al., 2016) and EuroSAT (Helber et al., 2019). These datasets are not typographic attack datasets, so we follow the procedure in (Azuma & Matsui, 2023a) to generate typographic attack images from these datasets. We also consider three more datasets designed for typographic attacks (results on table 2): Materzynska (Materzynska et al., 2022), PAINT (Ilharco et al., 2022), and RTA-100 (Azuma & Matsui, 2023a). Figure 8 shows some samples from the RTA-100 dataset. Additionally, for the Gender Bias results and analysis, we considered the VL-Bias Zhang et al. (2022a) dataset. Also, we consider Food101 (Bossard et al., 2014) for additional experiments, including saliency maps in section C.
Figure 8: Samples from RTA-100 dataset.
Table 3: Datasets statistics
@@
Appendix F Hyperparameter Sensitivity Analysis
@@
F.1 Effect of $\beta$
The hyper-parameter $\beta$ controls the deviation from the policy in the loss function 13. To assess its impact, we experiment with values between 0.01 and 1.5 on both the in-domain and zero-shot datasets.
@@ F.1 Effect of $\beta$
Refer to caption Refer to caption
(a) In-domain accuracy (b) Zero-shot accuracy
Figure 9: Ablation study on the effect of the hyper-parameter $\beta$ on classification accuracies for clean and typographic datasets. The results highlight the performance of DPO, IPO, and KTO models and their impact on in-domain accuracy and (b) zero-shot accuracy.
F.2 Ablation Study on the Regularization Loss $\mathcal{L}_{\text{reg}}$ and the Effect of $\lambda$
@@
Refer to caption Refer to caption
(a) In-domain accuracy (b) Zero-shot accuracy
Figure 10: Ablation study on the effect of the hyper-parameter $\lambda$ on classification accuracies for clean and typographic datasets. The results highlight the performance of DPO, IPO, and KTO models and their impact on in-domain accuracy and (b) zero-shot accuracy.
Appendix G Beta Moving Average vs. Exponential Moving Average
@@
7:     $l_{\text{reg}}\leftarrow\mathcal{L}_{\text{reg}}(\pi_{\theta_{t-1}},\pi_{\text{ref}};b_{\text{reg}})$triangleright Compute regulatory loss using $b_{\text{reg}}=\{(x,\tilde{x})\}$
@@
10:     Calculate $\alpha_{t}$ of the current model as in Eq. 35
11:     Update the BMA model $\theta^{\text{BMA}}_{t}$ as in Eq. 36
@@
Appendix H Sample Efficiency Analysis
In this section, we examine how efficiently our model learns with different dataset sizes. This analysis helps us understand how well the models perform with less data and how their accuracy changes as the dataset grows (see Figure 11). As expected, based on the foundation in Section 4.2 of (Azar et al., 2023a), when there is insufficient training data, DPO tends to overfit, leading to poor performance on both In-domain and Zero-shot datasets, while KTO performs better in this situation.
@@ Appendix H Sample Efficiency Analysis
Refer to caption Refer to caption
(a) In-domain accuracy (b) Zero-shot accuracy
Figure 11: Ablation study on the effect of dataset size on classification accuracies for clean and typographic datasets. The results highlight the performance of DPO, IPO, and KTO models and their impact on (a) in-domain accuracy and (b) zero-shot accuracy.
@@
H.1 Impact of Number of Typographic Instances per Image
@@
To investigate the effect of varying the number of typographic instances per image, we conducted experiments using 1, 3, 5, and 10 typographic variations for each image on the FOOD101 dataset (in-domain) and the SUN dataset (zero-shot). We evaluate the effect of varying the number of typographic instances per image on model performance. The results, summarized in Table 5, suggest that increasing the number of instances does not significantly improve accuracy, indicating that this augmentation technique may not be effective for enhancing the model’s generalization.

### v3

@@ Appendix B Proofs
$=\mathbb{E}_{(x_{1},y_{1},y_{2})\sim\mathcal{D}}\left[\left(\log\left(\frac{x_{1}}{x_{2}}\right)-\frac{\beta^{-1}}{2}\right)^{2}\right].$ (29)
@@
$=2\left(\log\left(\frac{x_{1}}{x_{2}}\right)-\frac{\beta^{-1}}{2}\right)\cdot\frac{1}{x_{1}}.$
@@ B.3 The Gradient of the DPO and IPO Objectives
Appendix C Features learned by our method
In Figure 6, we show saliency maps of vanilla CLIP and our fine-tuned model on images with typographic attacks. CLIP focuses on irrelevant parts of images, especially typographic texts, which results in false predictions. On the other hand, our model is capable of ignoring the text and focuses on the visual concept of the image, which results in a correct prediction.
ImageSaliency map (CLIP)Saliency map (fine-tuned)Refer to caption
Figure 6: Saliency maps of vanilla CLIP and our fine-tuned model given four different images with typographic attacks.
Further, in Figure 7, we provide the predictions of several models on another example image with a typographic attack. While the fine-tuned model remains robust, correctly classifying the attacked image, the original model misclassifies it. The inverse model confidently selects the typographic label but preserves classification accuracy.
(a) Example image showing dumplings after a typographic attack with misleading text (“clam chowder“).
a photo of a dumplinga photo of a clam chowderProbabilityInverse (typo.)Inverse (clean)Pretrained (typo)Pretrained (clean)Fine-tuned (typo.)Fine-tuned (clean)
(b) Predictions of six models.
Figure 7: Model predictions given an example image with a typographic attack.
These analyses showcase how applying our method forces the model to focus on the suitable features of the data rather than different sorts of attacks such as typographic.
@@ Appendix D Latent visualization using VQGAN
In this section, we further investigate our fine-tuned CLIP on typographical attack datasets. We use VQGAN-CLIP (Crowson et al., 2022) as the generative model which uses CLIP in its backbone. The results are shown in Figure 8. As can be seen, our fine-tuned CLIP completely ignores text in its image encoder representation and only focuses on visual concepts which increase its reliability and trustworthiness.
@@
Figure 8: Retrieved images using VQGAN-CLIP (Crowson et al., 2022) using the captions ”focus”, ”Love”, ”Male police” and ”Time” for image generation. As can be seen, vanilla CLIP puts the texts on the images, but our model fine-tuned on the typographical attack dataset SUN ignores such image texts and generates only visual concepts.
Appendix E Near Orthogonality of $W$
Previous works such as (Materzynska et al., 2022) also utilized a linear projection for their optimization objective. However, when these projections stray away from near orthogonal projections, our pretrained knowledge might be lost; this is a simple consequence of the following fact:
$\left.\begin{array}[]{ll}\tilde{f}(y,x)=\mathcal{I}(x)^{\top}W^{\top}W\mathcal{T}(y)\\ f(y,x)=\mathcal{I}(x)^{\top}\mathcal{T}(y)\end{array}\right\}\Rightarrow\tilde{f}(y,x)-f(y,x)=\mathcal{I}(x)^{\top}(W^{\top}W-I)\mathcal{T}(y)\\ \leq\|\mathcal{I}(x)\|\|\mathcal{T}(y)\|\|W^{\top}W-I\|\leq\|W^{\top}W-I\|,$ (35)
where the second equality is a result of CLIPs’ normalized embeddings, as it is evident from Eq. (35), our deviation from the base model is controlled by $\|W^{\top}W-I\|$, and ideally, we do not want the deviation to be too large. Previous works such as (Materzynska et al., 2022) added explicit loss objectives such as $R(W)=\|W^{\top}W-I\|^{F}$ to maintain proximity, and they have demonstrated the importance of this orthogonality loss. However, our method does not let $\pi$ and $\pi_{\text{ref}}$ diverge easily. Therefore, we expect unitary by default. After training the model using our method, considering that pretrained knowledge has been maintained, we already expect the matrix $W$ to be almost orthogonal. Nonetheless, we quantify this proximity in Figure 9.
$W^{\top}W$Deviation from Identity matrixSingular value distributionSingular value
(a) Model trained on achieving typographic attack robustness.
$W^{\top}W$Deviation from Identity matrixSingular value distributionSingular value
(b) Model trained on reversing gender understanding.
Figure 9: On the orthogonality of $W$: The left plots show the overall heatmaps of $W^{\top}W$, where it is evident that the matrix is close to the identity matrix. The middle plots display the residuals, which range between $-0.03$ and $0.03$, much smaller than the diagonal values. On the right, we show the distributions of the singular values, which are concentrated around 1.0, indicating that $W$ is close to being orthogonal and near the identity matrix.
The original CLIP model outputs normalized embeddings. Using an arbitrary linear projection can change the norm of the embeddings and hurt model performance and robustness. However, in our typical training scenario, this is not much of an issue since our transformation is close to an orthogonal transformation.
$\|Wx\|^{2}_{2}=x^{\top}W^{\top}Wx\leq\|x\|^{2}_{2}\|W^{\top}W\|=\|W^{\top}W\|=\|W^{\top}W-I+I\|\\ \leq\|W^{\top}W-I\|+\|I\|=1+\|W^{\top}W-I\|\approx 1\Rightarrow\frac{Wx}{\|Wx\|}\approx Wx$ (36)
However, when we amplify the singular values using matrix exponents, as discussed in Section 3.4, this difference grows exponentially; therefore, we should be careful and apply normalization.
Appendix F Datasets & Baselines
While human-curated preference sets are ideal, our methodology can still be effectively applied by designing the pretraining task in a semi-supervised manner. In both of our experiments, we did not rely on publicly available preference sets or human-curated preferences. Instead, we designed them ourselves, tailored to the specific task we wanted to fine-tune on. For instance, in the first experiment, we generated synthetic typographic attacks and used the original and targeted labels as preference labels. For debiasing the model, we used a dataset of images showing individuals of each gender performing various tasks and fine-tuned the model on an auxiliary task of reversing the model’s gender understanding, again without curated preferences. This demonstrates that preference optimization can be effectively applied using task-specific, semi-supervised strategies.
F.1 Datasets
@@
Table 3 presents the statistics of the datasets used in this paper. To evaluate the classification accuracy of our method on both original and typographic images (results in Table 1), we consider $9$ datasets: ImageNet-100 (Deng et al., 2009), Caltech101 (Fei-Fei et al., 2004), OxfordPets (Parkhi et al., 2012), StanfordCars (Krause et al., 2013), Flowers102 (Nilsback & Zisserman, 2008), FGVCAirCrafts (Maji et al., 2013), DTD (Cimpoi et al., 2014), SUN397 (Xiao et al., 2016) and EuroSAT (Helber et al., 2019). These datasets are not typographic attack datasets, so we follow the procedure in (Azuma & Matsui, 2023) to generate typographic attack images from these datasets. We also consider three real-world datasets designed for typographic attacks in Table 2: Materzynska (Materzynska et al., 2022), PAINT (Ilharco et al., 2022), and RTA-100 (Azuma & Matsui, 2023). Figure 10 shows some samples from the RTA-100 dataset. Additionally, for the Gender Bias results and analysis, we considered the VL-Bias (Zhang et al., 2022) dataset. Also, we consider Food101 (Bossard et al., 2014) for additional experiments, including saliency maps in Section C.
Table 2: Classification accuracy on real-world typographic datasets.
Figure 10: Samples of the RTA-100 dataset.
Table 3: Datasets statistics.
@@
F.2 Baselines
In the typographic attack experiments, we used the following works as our baseline.
  • (Materzynska et al., 2022). focus on disentangling the representation of written words and visual concepts in CLIP’s image encoder. They achieve this through orthogonal projections, creating subspaces that isolate or eliminate the model’s ability to process text. This disentanglement helps reduce text artifacts in image generation and offers defense against typographic attacks.
  • (Ilharco et al., 2022) introduce PAINT (Patching with Interpolation), a method that linearly interpolates between the weights of a model before and after fine-tuning on a specific patching task $(\theta_{\text{patch}}=\alpha\theta_{\text{zs}}+(1-\alpha)\theta_{\text{ft}})$. This approach improves accuracy on tasks where the zero-shot CLIP model performs poorly, while preserving performance on tasks like ImageNet. PAINT is computationally efficient, as it doesn’t require additional parameters or retraining from scratch. It also exhibits ”broad transfer,” where patching on one task can improve accuracy on related tasks, even with different classes.
  • (Azuma & Matsui, 2023) present Defense-Prefix (DP), a technique that inserts a DP token before a class name in text prompts to make it resilient to typographic attacks. This method focuses on modifying the text input to CLIP rather than changing the model itself. DP significantly enhances accuracy on typographic attack datasets while maintaining the model’s zero-shot capabilities. The authors also demonstrate its applicability to downstream tasks like object detection, where it effectively reduces the impact of typographic attacks without requiring additional training.
Table 1 does not include a comparison of our method with other baselines on the ImageNet dataset. This is because the experimental results of other methods were conducted on the ImageNetV2 validation subset, which was not publicly available at the time. Instead, we present results on a different subset of the ImageNetV2 dataset, referred to as the “matched-frequency“ subset (Recht et al., 2019). Table 4 provides a comparison of our method with the pretrained CLIP model under the typographic attack scenario.
Table 4: Classification accuracy results on the ImageNetV2 dataset.
Method O T CLIP 54.21 33.23 Ours (DPO) 53.80 49.18 Ours (IPO) 47.61 42.75 Ours (KTO) 53.17 50.35
Appendix G Hyperparameter Sensitivity Analysis
@@
G.1 Effect of $\beta$
The hyper-parameter $\beta$ controls the deviation from the policy in the loss function in Eq. (14). To assess its impact, we experiment with values between 0.01 and 1.5 on both the in-domain and zero-shot datasets.
@@ G.1 Effect of $\beta$
(a) In-domain accuracy.
(b) Zero-shot accuracy.
Figure 11: Ablation study on the effect of the hyper-parameter $\beta$ on classification accuracies for clean and typographic datasets. The results highlight the performance of DPO, IPO, and KTO models and their impact on in-domain and zero-shot accuracy.
G.2 Ablation Study on the Regularization Loss $\mathcal{L}_{\text{reg}}$ and the Effect of $\lambda$
@@
(a) In-domain accuracy.
(b) Zero-shot accuracy.
Figure 12: Ablation study on the effect of the hyper-parameter $\lambda$ on classification accuracies for clean and typographic datasets. The results highlight the performance of DPO, IPO, and KTO models and their impact in-domain and zero-shot accuracy.
Appendix H Beta Moving Average vs. Exponential Moving Average
@@
7:     $l_{\text{reg}}\leftarrow\mathcal{L}_{\text{reg}}(\pi_{\theta_{t-1}},\pi_{\text{ref}};b_{\text{reg}})$triangleright Compute regularization loss using $b_{\text{reg}}=\{(x,\tilde{x})\}$
@@
10:     Calculate $\alpha_{t}$ of the current model as in Eq. (38)
11:     Update the BMA model $\theta^{\text{BMA}}_{t}$ as in Eq. (39)
@@
Appendix I Sample Efficiency Analysis
In this section, we examine how efficiently our model learns with different dataset sizes. This analysis helps us understand how well the models perform with less data and how their accuracy changes as the dataset grows (see Figure 13). As expected, based on the foundation in Section 4.2 of (Gheshlaghi Azar et al., 2024), when there is insufficient training data, DPO tends to overfit, leading to poor performance on both In-domain and Zero-shot datasets, while KTO performs better in this situation.
@@ Appendix I Sample Efficiency Analysis
(a) In-domain accuracy.
(b) Zero-shot accuracy.
Figure 13: Ablation study on the effect of dataset size on classification accuracies for clean and typographic datasets. The results highlight the performance of DPO, IPO, and KTO models and their impact on in-domain and zero-shot accuracy.
@@
I.1 Impact of Number of Typographic Instances per Image
@@
To investigate the effect of varying the number of typographic instances per image, we conducted experiments using 1, 3, 5, and 10 typographic variations for each image on the FOOD101 dataset (in-domain) and the SUN dataset (zero-shot). We evaluate the effect of varying the number of typographic instances per image on model performance. The results, summarized in Table 6, suggest that increasing the number of instances does not significantly improve accuracy, indicating that this augmentation technique may not be effective for enhancing the model’s generalization further.
Appendix J Performance Comparison of DPO, IPO, and KTO
In this section, we provide a detailed analysis of the performance of DPO, IPO, and KTO in our framework. The comparison focuses on key aspects such as sample efficiency, sensitivity to the $\beta$ hyperparameter, and performance across different datasets.
J.1 Sample Efficiency
As discussed in Section I and illustrated in Figure 13, KTO demonstrates superior performance in scenarios where training data is scarce. This result aligns with our expectations, as KTO transforms each preference triplet $(x,y_{w},y_{l})\in\mathcal{D_{\text{pref}}}$ into two samples: $(x,y_{\text{desired}})$ and $(x,y_{\text{undesired}})$. By effectively doubling the number of training samples, KTO achieves greater sample efficiency, making it particularly beneficial in data-constrained settings.
In contrast, DPO performs the poorest among the three methods when training data is insufficient. This is consistent with the findings of (Gheshlaghi Azar et al., 2024), where DPO’s tendency to overfit in low-data regimes is highlighted. Overfitting leads to poor generalization, resulting in suboptimal performance on both in-domain and zero-shot datasets. IPO, while not as sample-efficient as KTO, outperforms DPO in such scenarios due to its relatively balanced approach to learning from preference data.
J.2 Sensitivity to $\beta$
Figure 11 highlights the sensitivity of each method to the $\beta$ hyperparameter, which governs the strength of the policy regularization. IPO is particularly sensitive to variations in $\beta$. Large $\beta$ values lead to a uniform policy distribution $\pi_{\theta}$, causing the model to assign significant probabilities to all actions. This behavior reduces specificity and harms performance, particularly on zero-shot datasets.
For both DPO and KTO, excessively high $\beta$ values similarly degrade performance. However, KTO exhibits slightly greater robustness compared to DPO in this regard, likely due to its ability to leverage additional training samples. To achieve reasonable performance with IPO, stronger constraints are required to retain the policy close to the reference model, as noted by (Gheshlaghi Azar et al., 2024). Fine-tuning $\beta$ is therefore crucial for optimizing IPO performance.
J.3 Performance Across Different Datasets
Our ablation results in Figure 13 and the reported results in Table 1 indicate that with sufficient training data, all three methods—DPO, IPO, and KTO—exhibit comparable performance on both original and adversarial datasets. This convergence in performance is expected, as the final loss function in all three methods is derived from the same main objective (Eq. (3)), albeit with different design principles and optimization approaches. However, under specific conditions, their distinct characteristics lead to notable differences, highlighting the unique strengths and limitations of each method.
Nonetheless, specific strengths emerge in certain contexts:
  • Synthetic Datasets: As shown in Table 1, KTO outperforms DPO and IPO on synthetic datasets, achieving better average performance on both the original and typographic variants.
  • Real-World Datasets: As shown in Table 2, IPO achieves the best average performance among the three methods on real-world datasets.
J.4 Weight Analysis of Different Variants
Regarding the specific analysis of the weight of different variants according to Eq. (10), we provide the following explanation:
  • DPO: The gradient weight is given by:
    $w_{\text{pref}}(y_{w},y_{l};x)=\sigma(-\beta h_{\pi_{\theta}}(y_{w},y_{l},x)),$
    where $\sigma(\cdot)$ is the sigmoid function. As $-\beta h_{\pi_{\theta}}(y_{w},y_{l},x)$ increases, the weight decreases, which results in lower emphasis on inputs with higher values of $h_{\pi_{\theta}}$, as they are already good enough. This deemphasizing prevents the model from straying too far from the base model.
  • IPO: The gradient weight is defined as:
    $w_{\text{pref}}(y_{w},y_{l};x)=\left(\frac{1}{2\beta}-h_{\pi_{\theta}}(y_{w},y_{l},x)\right).$
    Here, as $h_{\pi_{\theta}}(y_{w},y_{l},x)$ increases, the weight linearly decreases, becoming zero at $h_{\pi_{\theta}}(y_{w},y_{l},x)=\frac{1}{2\beta}$. In effect, this would stop the model from straying further than $\frac{1}{2\beta}$, ideally optimizing to around $h_{\pi_{\theta}}(y_{w},y_{l},x)\approx\frac{1}{2\beta}$.
  • Other Optimization Schemes: Any other preference optimization scheme that uses the differential reward can be studied. One such case could be ROPO, as introduced by (Liang et al., 2024b).
We compare the different weighting schemes in Figure 14.
Figure 14: Visualization of different weighting schemes.
Figure 15: The violin plot represents the distribution of the $h$ function in DPO method across different datasets. From left to right, the datasets progressively become less challenging. By ’challenging,’ we refer to the performance of the pretrained model on the typographic variation of each dataset. As seen, the dataset on the far right corresponds to our training set, which shows higher and more accurate $h$ values.
Figure 15 illustrates how the $h$ function correlates with dataset difficulty, highlighting the relationship between model performance and the inherent challenges posed by different datasets.
Appendix K Experiments on Gradient-based Attacks
In this experiment, we evaluate the effectiveness of our proposed method in enhancing model robustness against gradient-based attacks, specifically the Projected Gradient Descent (PGD) attack. PGD is a widely used method for generating adversarial examples by iteratively perturbing the input to maximize the model’s loss.
The setup is as follows: we use targeted PGD attacks to create an image $x^{\prime}$, which the model predicts the caption $y_{l}$ for instead of $y_{w}$, where $y_{l}$ is the target caption of the adversary. Our methodology from the typographic attack section remains largely the same, with a few adjustments. Specifically, we attack the model in an online fashion during training. The training accuracies of the model on clean and perturbed data are illustrated in Figures 16 and 17. Additionally, our results on the validation set of the dataset are provided in Table 7.
In these experiments, for $\epsilon=4/255$, we used 700 iterations with a batch size of 256, and for $\epsilon=2/255$, we used 1000 iterations with a batch size of 256.
Overall, it is evident that the clean data accuracy remains the same while adversarial robustness is increasing.
(a) Adversarially perturbed data accuracy
(b) Clean data accuracy
Figure 16: Results of training on adversarial robustness, CIFAR10 dataset, $\epsilon=2/255$.
(a) Adversarially perturbed data accuracy
(b) Clean data accuracy
Figure 17: Results of training on adversarial robustness, CIFAR10 dataset, $\epsilon=4/255$.
Dataset Epsilon ($\epsilon$) Accuracy on Adversarial Images Accuracy on Clean Images
CIFAR-10 $2/255$ 62.89% 98.44%
CIFAR-10 $4/255$ 21.88% 97.27%
Table 7: Evaluation of Model Accuracy on CIFAR-10 Dataset for Clean and Adversarially Perturbed Images.